# Rate-Limited Closed-Loop Distributed ISAC Systems: An Autoencoder Approach

### v1

@@ I Introduction
Recent advances in 5G/6G networks have enabled closed-loop distributed integrated sensing and communication (ISAC) systems [1], where sensing, communication, and control are tightly coupled in real time [2]. These systems show strong potential in applications such as autonomous driving [3], industrial automation [4], and UAVs [5]. Unlike conventional decoupled designs, closed-loop ISAC jointly optimizes sensing, communication, and control to support robust estimation and decision-making under rate-limited conditions. This integration improves efficiency and resilience, especially under bandwidth and network dynamics.
@@ II Closed-loop Distributed ISAC System Model
Closed-loop distributed ISAC systems, as depicted in Fig. 1, tightly integrate sensing, communication, and control, and have broad applications in domains such as autonomous driving, industrial automation, and drone coordination. In this section, we outline a general framework for such systems.
@@
  • Sensor Model: The environment is equipped with $S$ sensors, (e.g., sensing base stations, cameras, LiDARs) that continuously sense and track the controllable objects. These sensors collect high-dimensional raw observations (such as channel measurements, images, or dense point clouds) and forward them to a fusion center for state estimation and prediction. We denote the observation from sensor $s\in\{1,\ldots,S\}$ by $\bm{y}_{t}^{s}\sim p_{\text{meas}}^{s}(\bm{y}_{t}^{s}|\bm{x}_{t})$, where $p_{\text{meas}}^{s}(\cdot|\bm{x}_{t})$ is the observation model of sensor $s$.
  • Digital Twin and State Tracker: The fusion center maintains a real-time model of the tracked objects in the form of a digital twin (DT). Given observations $\mathcal{Y}_{t}=\bm{y}_{t}^{1:S}$ and controls $\mathcal{U}=\bm{u}_{1:t}$, the DT models the posterior $q(\bm{x}_{t}|\mathcal{Y}_{t},\mathcal{U}_{t-1})$, and the kinematics $p_{\text{DT}}(\bm{x}_{t+k}|\bm{x}_{t},\bm{u}_{t})$, for any $k>0$. Note that $p_{\text{DT}}(\bm{x}_{t+k}|\bm{x}_{t},\bm{u}_{t})$ may or may not be identical to $p_{\text{obj}}(\bm{x}_{t+k}|\bm{x}_{t},\bm{u}_{t})$, depending on the fidelity of the DT. The DT then provides a prediction on the states of the tracked objects. When measurements $\mathcal{Y}_{t+1}$ are provided, with associated measurement model $p_{\text{DT}}(\mathcal{Y}_{t+1}|\bm{x}_{t+1})$, which under ideal conditions is $p_{\text{DT}}(\mathcal{Y}_{t+1}|\bm{x}_{t+1})=\prod_{s=1}^{S}p_{\text{meas}}^{s}(\bm{y}_{t+1}^{s}|\bm{x}_{t+1})$, the state tracker updates the posterior as
@@
  • Controller: Assuming a separation between state tracking and control, the controller takes as input the posterior distribution from the DT and determines the control command: $\bm{u}_{t+1}\sim\pi(q(\bm{x}_{t+1}|\mathcal{Y}_{t+1},\mathcal{U}_{t}))$, where $\pi(\cdot)$ denotes the control policy. This policy is determined a prior to optimize a certain control objective. Examples include established control methodologies such as model predictive control (MPC) or reinforcement learning (RL). These commands aim at specific operational goals, including guiding vehicles toward desired positions, or stabilizing drone flight trajectories. Finally, these control commands are transmitted to the controlled objects to influence their subsequent state transitions described by the conditional probability distribution $p_{\text{obj}}(\bm{x}_{t+2}|\bm{x}_{t+1},\bm{u}_{t+1})$ thus closing the feedback loop.
  • Rate-Limited Links: Given the significant communication constraints typical in distributed systems, directly transmitting these raw observations $\bm{y}_{t}^{1:S}$ could rapidly exceed the available network bandwidth [1, 17, 18]. Therefore, sensors must apply local data compression methods and resource allocation schemes, extracting and transmitting only task-relevant information, thus maintaining accurate state estimation at the fusion center without violating bandwidth limitations. Semantic and goal-oriented communication strategies have emerged as promising solutions to enhance data compression efficiency by capturing the most relevant information directly related to the control objectives.
@@ III-A Rate-Limited Problem Description
As bandwidth is limited in closed-loop distributed ISAC systems, sensors locally compress their observations and collaboratively allocate resources to ensure that critical information reaches the fusion center. Specifically, each sensor $s$ encodes its raw observation $\bm{y}_{t}^{s}$ into a compressed representation $\bm{z}_{t}^{s}=\mathcal{F}_{\text{comp}}^{s}(\bm{y}_{t}^{s};\bm{\theta}_{\text{comp}}^{s})$ using a local compression function $\mathcal{F}_{\text{comp}}^{s}$, parameterized by $\bm{\theta}_{\text{comp}}^{s}$. The compressed data is transmitted over a rate-limited link to the fusion center, where it is reconstructed via a corresponding decompression function $\mathcal{F}_{\text{rec}}^{s}$ with parameters $\bm{\theta}_{\text{rec}}^{s}$. The reconstruction process is defined as $\hat{\bm{y}}_{t}^{s}=\mathcal{F}_{\text{rec}}^{s}(\bm{z}_{t}^{s};\bm{\theta}_{\text{rec}}^{s})$. To meet communication constraints, the total transmission dimension from all sensors at each time step must satisfy $\sum_{s=1}^{S}\dim(\bm{z}_{t}^{s})\leq r_{t},$ where $\dim(\bm{z}_{t}^{s})$ denotes the compression dimension for sensor $s$. Generally, both the transmission dimension budget $r_{t}$ and the compression dimension $\bm{z}_{t}^{s}$ may vary over time, enabling dynamic adaptation to channel conditions, task requirements, or estimation uncertainty.
III-B Three Levels of Error in DISAC
To characterize system performance under rate-limited communication, we consider three levels of error, each corresponding to a different stage in the closed-loop ISAC pipeline:
  • Reconstruction Error: Measures the difference between the original and reconstructed sensor observations, indicating how effectively the AE compresses and reconstructs the data.
  • Estimation Error: Quantifies the accuracy of the state estimation by measuring the deviation between the fusion center’s estimated state and the true system state, reflecting the direct effect of compression fidelity on state tracking performance.
  • Control Error: Captures the ultimate impact of compression on system objectives by evaluating the incurred control cost within the closed-loop operation, directly linking compression quality to control effectiveness.
@@ III-C Autoencoder-Based Compression Method
In this study, we focus on designing compression methods that explicitly optimize for reconstruction accuracy. [^1] While our current framework focuses on reconstruction, future work will incorporate task-driven losses directly tied to estimation or control objectives. We focus on how minimizing reconstruction error impacts state estimation and control performance in closed-loop distributed ISAC systems. To this end, we propose an AE-based compression framework. An AE is a neural network architecture composed of an encoder and a decoder [19]. The encoder compresses the input observation into a low-dimensional latent representation. The decoder attempts to reconstruct the original observation from this compressed data.
@@ IV Case Study: Closed-Loop LQR
To investigate the proposed compression framework, we study a closed-loop LQR problem [13], which serves as a fundamental instance of the distributed ISAC setting, where optimal state estimators and controllers are known and available in closed form. Our goals are: (i) to analyze how system performance depends on the interplay among state, observation, and compression dimensions; and (ii) to study how to allocate limited communication resources across sensors with different noise levels to enhance control performance.
@@ IV-A Scenario Description
To gain practical insights from the general framework of closed-loop ISAC system, we consider an LQR system, as shown in Fig. 2. Sensors observe a plant and transmit observations over rate-limited links to a fusion center, where a state tracker uses a Kalman filter to estimate the plant’s state. The controller then generates control commands based on the estimation, forming a closed loop.
@@ IV-A1 State Model
The system dynamics are given by
$\bm{x}_{t+1}=\bm{A}\bm{x}_{t}+\bm{B}\bm{u}_{t}+\bm{v}_{t},$ (3)
where $\bm{x}_{t}\in\mathbb{R}^{N_{x}}$ is the state of dimension $N_{x}$, and $\bm{u}_{t}\in\mathbb{R}^{N_{u}}$ is the control of dimension $N_{u}$. The matrices $\bm{A}\in\mathbb{R}^{N_{x}\times N_{x}}$ and $\bm{B}\in\mathbb{R}^{N_{x}\times N_{u}}$ represent the known system dynamics and control input mappings, respectively. The process noise $\bm{v}_{t}\sim\mathcal{N}(0,\bm{Q})$ is with covariance matrix $\bm{Q}\in\mathbb{R}^{N_{x}\times N_{x}}$. At each time step, sensor $s$ provides linear measurements:
@@ IV-A2 State Tracker
Prior to the state tracker, the compressed measurements are decompressed using $\hat{\bm{y}}_{t}^{s}=\mathcal{F}_{\text{rec}}^{s}(\bm{z}_{t}^{s};\bm{\theta}_{\text{rec}}^{s})$, and then provided to the Kalman filter. The Kalman filter combines them with the system model to estimate $\hat{\bm{x}}_{t|t}$ and its associated covariance matrix $\bm{\Sigma}_{t|t}$, following standard prediction and update steps [20]. The prediction step relies on the DT of the real system (3). When the observation $\bm{y}_{t}^{s}$ is perfectly reconstructed and the DT models the system correctly, the Kalman filter is optimal. However, when there is reconstruction error in $\hat{\bm{y}}_{t}^{s}$, the Kalman filter becomes suboptimal.
@@ IV-A3 Controller
In the controller, the LQR framework is used to compute the control input $\bm{u}_{t}$ that minimizes the infinite-horizon cost
@@
where $\tilde{\bm{x}}_{t}=\bm{x}_{t}-\bm{x}_{\mathrm{desired}}$ denotes the deviation from the target state, and $\bm{Q}_{\mathrm{goal}}$ and $\bm{R}_{\mathrm{goal}}$ are matrices penalizing state deviation and control effort, respectively. When the observation $\bm{y}$ is perfectly, the optimal control input follows the LQR policy:
$\bm{u}_{t}=-\bm{K}\,\hat{\bm{x}}_{t|t},$ (6)
where $\bm{K}$ can be computed as described in [21].
It is important to note that without compression and reconstruction, the LQR system operates at its optimum. However, compression over rate-limited links introduces errors that may destabilize the system. In our experiments, we fix the LQR system parameters so that the system operates stably with a stationary data distribution when the reconstruction error is maintained below a critical threshold. This setup allows us to focus on the static effects of compression dimensions and resource allocation strategies on system performance, without considering their dynamic variations.
@@ IV-B2 AE Implementation
In our implementation, both the encoder and decoder are designed as fully-connected multi-layer perceptrons. The encoder consists of four linear layers with ReLU activations, with the number of neurons in each layer set to 512, 1024, 512, and the target compression dimension $\dim(\bm{z}_{t}^{s})$, respectively. The decoder mirrors this structure symmetrically to reconstruct the observation from the latent representation. We run 1000 rounds of 200 steps each to collect 200,000 training samples, then test on another 1000 rounds (200 steps each).
@@ IV-C2 Lossless Compression at the State Dimension
We plot the results for the case where the compression dimension is fixed to the system’s state dimension, i.e., $\dim(\bm{z}_{t}^{s})=N_{x}=4$ in Fig. 4. As shown in Fig. 4LABEL:sub@fig:MSE_CompDim4_diffObs, the reconstruction error is nearly identical in both online and offline settings for PCA and AE. As the observation dimension increases, reconstruction error also increases due to a higher compression ratio, which leads to more information loss. However, as shown in Fig. 4LABEL:sub@fig:state_CompDim4_diffObs and Fig. 4LABEL:sub@fig:LQR_CompDim4_diffObs, both PCA and AE achieve state estimation and LQR performance comparable to the uncompressed baseline, indicating lossless at this compression level. Furthermore, the state estimation error and LQR cost decrease as the observation dimension increases, showing the benefit of more sensing information.
@@ IV-D1 Performance with Fixed Transmission Dimension
For the multi-sensor scenario, we consider a system with two distributed sensors, each having an observation dimension of 20. sensor 1 has low noise with $\bm{R}^{1}=\bm{I}$, while sensor 2 has higher noise with $\bm{R}^{2}=10\bm{I}$. The resource allocation strategy is defined as a pair $(\dim(\bm{z}_{t}^{1}),\dim(\bm{z}_{t}^{2}))$, representing the compression dimensions allocated to sensor 1 and sensor 2, respectively. We first evaluate system performance under a fixed total transmission dimension $r_{t}$, comparing different allocation strategies. We then examine how the optimal performance and corresponding allocation under different $r_{t}$.
@@
Fig. 6 illustrates the system performance when the total transmission dimension is fixed at 6, i.e., $\dim(\bm{z}_{t}^{1})+\dim(\bm{z}_{t}^{2})=r_{t}=6$. As shown in Fig. 6LABEL:sub@fig_fixed_transmission_dimension_LQR, when the compression dimension of sensor 1 increases, the LQR cost and state estimation error first decrease and then increase. The best result occurs at the allocation (4,2). Since sensor 1 (the low-noise sensor) provides better observations for state estimation, as its compression dimension increases, performance improves until lossless compression is effectively achieved at dimension 4. Beyond that, allocating additional resources to sensor 1 (e.g., (5,1)) reduces the transmission dimension available to sensor 2, discarding potentially valuable complementary information. Meanwhile, Fig. 6LABEL:sub@fig_fixed_transmission_dimension_MSE shows the reconstruction error of sensors 1 and 2 across allocation strategies. As sensor 1’s compression dimension increases, its reconstruction MSE decreases consistently. For sensor 2, as the compression dimension of sensor 1 increases, the MSE first decreases and then increases. This rise is due to two factors: Initially, when the compression dimension of sensor 1 is low but that of sensor 2 is high, the LQR system deviates from its stable state, causing the sensor observations to fall outside the training distribution of AE and degrading the reconstruction quality. Later, as the compression dimension of sensor 1 increases, sensor 2 suffers from lossy compression at a lower compression dimension, and the reconstruction error increases again.
@@
Figure 7: Performance under different total transmission dimensions.

### v2

# Observation Compression in Rate-Limited Closed-Loop Distributed ISAC Systems: From Signal Reconstruction to Control

@@ I Introduction
Recent advances in 5G/6G networks have enabled closed-loop distributed integrated sensing and communication (ISAC) systems [1], where sensing, communication, and control are tightly coupled in real time [2]. These systems show strong potential in applications such as autonomous driving [3], industrial automation [4], and UAVs [5]. Unlike conventional decoupled designs, closed-loop ISAC jointly optimizes sensing, communication, and control for robust estimation and decision-making under rate limits, enhancing efficiency and resilience amid bandwidth and network dynamics..
@@ II Closed-loop Distributed ISAC System Model
Closed-loop distributed ISAC systems, as depicted in Fig. 1, integrate sensing, communication, and control, with applications in autonomous driving and UAV coordination. This section outlines a general framework for such systems.
@@
  • Sensor Model: The environment is equipped with $S$ sensors, (e.g., full-duplex ISAC-enabled base stations [15]) that continuously sense and track the controllable objects. These sensors collect high-dimensional raw observations (such as channel measurements, images, or dense point clouds) and forward them to a fusion center for state estimation and prediction. We denote the observation from sensor $s\in\{1,\ldots,S\}$ by $\bm{y}_{t}^{s}\sim p_{\text{meas}}^{s}(\bm{y}_{t}^{s}|\bm{x}_{t})$, where $p_{\text{meas}}^{s}(\cdot|\bm{x}_{t})$ is the observation model of sensor $s$.
  • Digital Twin and State Tracker: The fusion center maintains a real-time model of the tracked objects in the form of a digital twin (DT). Given observations $\mathcal{Y}_{t}=\bm{y}_{t}^{1:S}$ and controls $\mathcal{U}=\bm{u}_{1:t}$, the DT models the posterior $q(\bm{x}_{t}|\mathcal{Y}_{t},\mathcal{U}_{t-1})$, and the kinematics $p_{\text{DT}}(\bm{x}_{t+k}|\bm{x}_{t},\bm{u}_{t})$, for any $k>0$. Note that $p_{\text{DT}}(\bm{x}_{t+k}|\bm{x}_{t},\bm{u}_{t})$ may or may not be identical to $p_{\text{obj}}(\bm{x}_{t+k}|\bm{x}_{t},\bm{u}_{t})$, depending on the fidelity of the DT. The DT then provides a prediction on the states of the tracked objects. Given observations $\mathcal{Y}{t+1}$ with measurement model $p_{\text{DT}}(\mathcal{Y}_{t+1}|\bm{x}_{t+1})=\prod_{s=1}^{S}p_{\text{meas}}^{s}(\bm{y}_{t+1}^{s}|\bm{x}_{t+1})$, the state tracker updates the posterior as
@@
  • Controller: Assuming a separation between state tracking and control, the controller takes as input the posterior distribution from the DT and determines the control command: $\bm{u}_{t+1}\sim\pi(q(\bm{x}_{t+1}|\mathcal{Y}_{t+1},\mathcal{U}_{t}))$, where $\pi(\cdot)$ denotes the control policy. This policy is determined to optimize a control objective, using methods such as model predictive control (MPC) or reinforcement learning (RL). The resulting commands pursue goals like guiding vehicles to target positions or stabilizing UAV trajectories. These commands are then transmitted to the controlled objects, influencing subsequent state transitions via $p_{\text{obj}}(\bm{x}{t+2}|\bm{x}{t+1},\bm{u}_{t+1})$.
  • Rate-Limited Links: Given the significant communication constraints typical in distributed systems, directly transmitting these raw observations $\bm{y}_{t}^{1:S}$ could rapidly exceed the available network bandwidth [1, 18, 19]. Therefore, sensors must apply local data compression methods and resource allocation schemes, extracting and transmitting only task-relevant information, thus maintaining accurate state estimation at the fusion center without violating bandwidth limitations. Goal-oriented semantic communication strategies have emerged as promising solutions to enhance data compression efficiency by capturing the most relevant information directly related to the control objectives.
@@ III-A Rate-Limited Problem Description
As bandwidth is limited in closed-loop distributed ISAC systems, sensors locally compress their observations and collaboratively allocate resources to ensure that critical information reaches the fusion center. Specifically, each sensor $s$ encodes its raw observation $\bm{y}_{t}^{s}$ into a compressed representation $\bm{z}_{t}^{s}=\mathcal{F}_{\text{comp}}^{s}(\bm{y}_{t}^{s};\bm{\theta}_{\text{comp}}^{s})$ using a local compression function $\mathcal{F}_{\text{comp}}^{s}$, parameterized by $\bm{\theta}_{\text{comp}}^{s}$. The compressed data is transmitted over a rate-limited link [^1] For the sake of clarity, these links are assumed to be error-free and do not incur any latency. to the fusion center, where it is reconstructed via a corresponding decompression function $\mathcal{F}_{\text{rec}}^{s}$ with parameters $\bm{\theta}_{\text{rec}}^{s}$. The reconstruction process is defined as $\hat{\bm{y}}_{t}^{s}=\mathcal{F}_{\text{rec}}^{s}(\bm{z}_{t}^{s};\bm{\theta}_{\text{rec}}^{s})$. To meet communication constraints, the total transmission dimension from all sensors at each time step must satisfy $\sum_{s=1}^{S}\dim(\bm{z}_{t}^{s})\leq r_{t},$ where $\dim(\bm{z}_{t}^{s})$ denotes the compression dimension for sensor $s$. Generally, both the transmission dimension budget $r_{t}$ and the compression dimension $\bm{z}_{t}^{s}$ may vary over time, enabling dynamic adaptation to channel conditions, task requirements, or estimation uncertainty. [^2] While compressed observations reduce transmission delay and thereby reduce sensing latency, the effects of control frequency and sensing latency are not considered in this work [20].
III-B Three Levels of Error in Distributed ISAC
To characterize performance under rate-limited communication, we consider three error levels, each tied to a stage of the closed-loop ISAC pipeline:
  • L1 – Reconstruction Error: Measures the difference between the original and reconstructed observations, indicating the compression and reconstruction effectiveness.
  • L2 – Estimation Error: Quantifies the accuracy of the state estimation by measuring the deviation between the fusion center’s estimation and the true system state, reflecting the impact of compression fidelity on tracking performance.
  • L3 – Control Error: Captures the ultimate impact of compression on system objectives by evaluating the incurred control cost within the closed-loop operation, directly linking compression quality to control effectiveness.
@@ III-C Autoencoder-Based Compression Method
In this study, we focus on designing compression methods that explicitly optimize for L1 – reconstruction error. [^3] Although distributed sensors may cause interference or complementary gains, this work focuses on observation compression in closed-loop distributed ISAC systems and ignores such effects. Future work will incorporate task-driven losses tied to estimation or control objectives. We focus on how minimizing reconstruction error impacts state estimation and control performance in closed-loop distributed ISAC systems. To this end, we develop an AE-based compression framework tailored for minimizing reconstruction error. An AE is a neural network architecture composed of an encoder and a decoder [21]. The encoder compresses the input observation into a low-dimensional latent representation. The decoder attempts to reconstruct the original observation from this compressed data.
@@ IV Case Study: Closed-Loop LQR
To evaluate the proposed compression framework, we study a closed-loop LQR problem [13], a fundamental paradigm in control theory that provides closed-form solutions and strong theoretical guarantees, and serves as a representative instance of the distributed ISAC setting. It serves as a first-order approximation for many more complex nonlinear control problems and has been extensively used to study estimation and control under communication constraints [13]. The schematic diagram of the LQR system is shown in Figure 2. Our goals are: (i) to analyze how system performance depends on the interplay among state, observation, and compression dimensions; and (ii) to study how to allocate limited communication resources across sensors with different noise levels to enhance control performance.
@@ IV-A Scenario Description
In distributed LQR systems, sensors send rate-limited observations of the plant to a fusion center, where a Kalman filter estimates the state. The controller then generates commands from this estimate, forming the closed loop.
@@ IV-A1 State Model
The system dynamics are given by $\bm{x}_{t+1}=\bm{A}\bm{x}_{t}+\bm{B}\bm{u}_{t}+\bm{v}_{t}$, where $\bm{x}_{t}\in\mathbb{R}^{N_{x}}$ is the state of dimension $N_{x}$, and $\bm{u}_{t}\in\mathbb{R}^{N_{u}}$ is the control of dimension $N_{u}$. The matrices $\bm{A}\in\mathbb{R}^{N_{x}\times N_{x}}$ and $\bm{B}\in\mathbb{R}^{N_{x}\times N_{u}}$ represent the known system dynamics and control input mappings, respectively. The process noise $\bm{v}_{t}\sim\mathcal{N}(0,\bm{Q})$ is with covariance matrix $\bm{Q}\in\mathbb{R}^{N_{x}\times N_{x}}$. At each time step, sensor $s$ provides linear measurements:
@@ IV-A2 State Tracker
Prior to the state tracker, the compressed measurements are decompressed using $\hat{\bm{y}}_{t}^{s}=\mathcal{F}_{\text{rec}}^{s}(\bm{z}_{t}^{s};\bm{\theta}_{\text{rec}}^{s})$, and then provided to the Kalman filter. The Kalman filter combines them with the system model to estimate $\hat{\bm{x}}_{t|t}$ and its associated covariance matrix $\bm{\Sigma}_{t|t}$, following prediction and update steps [22]. The prediction step relies on the DT of the real system (II). When the observation $\bm{y}_{t}^{s}$ is perfectly reconstructed and the DT models the system correctly, the Kalman filter is optimal. However, when there is reconstruction error in $\hat{\bm{y}}_{t}^{s}$, the Kalman filter becomes suboptimal.
@@ IV-A3 Controller
In the controller, the LQR computes the control input $\bm{u}_{t}$ by minimizing the infinite-horizon cost:
@@
where $\tilde{\bm{x}}_{t}=\bm{x}_{t}-\bm{x}_{\mathrm{desired}}$ denotes the deviation from the target state, and $\bm{Q}_{\mathrm{goal}}$ and $\bm{R}_{\mathrm{goal}}$ are matrices penalizing state deviation and control effort, respectively. When the observation $\bm{y}$ is perfect, the optimal control input follows the LQR policy: $\bm{u}_{t}=-\bm{K}\,\hat{\bm{x}}_{t|t}$, where $\bm{K}$ can be computed as described in [23].
Without compression, the LQR system achieves optimal performance. However, compression over rate-limited links introduces errors that may destabilize the system. In our experiments, we fix the LQR parameters so that the system remains stable with a stationary data distribution when the reconstruction error is below a critical threshold. This setup enables us to focus on the static impact of compression dimensions and resource allocation strategies on performance, without dynamic variations.
@@ IV-B2 AE Implementation
In our implementation, both the encoder and decoder are designed as fully-connected multi-layer perceptrons. The encoder consists of four linear layers with ReLU activations, with the number of neurons in each layer set to 512, 1024, 512, and the target compression dimension $\dim(\bm{z}_{t}^{s})$, respectively. The decoder mirrors this structure symmetrically to reconstruct the observation from the latent representation. We run 1000 rounds of 200 steps each to collect 200,000 training samples, then test on another 1000 rounds (200 steps each). We compare the AE with classical PCA, expected to be optimal for this linear problem, aiming to show that the AE can at least match PCA’s performance.
@@ IV-C2 Lossless Compression at the State Dimension
We plot the results for the case where the compression dimension is fixed to the system’s state dimension, i.e., $\dim(\bm{z}_{t}^{s})=N_{x}=4$ in Fig. 4. As shown in Fig. 4LABEL:sub@fig:MSE_CompDim4_diffObs, the reconstruction error is nearly identical in both online and offline settings for PCA and AE. As the observation dimension increases, reconstruction error also increases due to a higher compression ratio, which leads to more information loss. However, as shown in Fig. 4LABEL:sub@fig:state_CompDim4_diffObs and Fig. 4LABEL:sub@fig:LQR_CompDim4_diffObs, both PCA and AE achieve state estimation and LQR performance comparable to the uncompressed baseline, indicating lossless at this compression level. Furthermore, both state estimation error and LQR cost decrease with higher observation dimensions, highlighting the benefit of richer sensing information.
@@ IV-D1 Performance with Fixed Transmission Dimension
In the multi-sensor scenario, we consider two distributed sensors, each with an observation dimension of 20. Sensor 1 has low noise with $\bm{R}^{1}=\bm{I}$, while sensor 2 has higher noise with $\bm{R}^{2}=10\bm{I}$. The resource allocation strategy is specified as $(\dim(\bm{z}_{t}^{1}),\dim(\bm{z}_{t}^{2}))$, denoting the compression dimensions assigned to sensors 1 and 2, respectively. We first evaluate performance under a fixed total transmission dimension $r_{t}$, comparing different allocation strategies. We then examine the optimal performance and corresponding allocations across varying $r_{t}$.
@@
Fig. 6 illustrates the system performance when the total transmission dimension is fixed at 6, i.e., $\dim(\bm{z}_{t}^{1})+\dim(\bm{z}_{t}^{2})=r_{t}=6$. As shown in Fig. 6LABEL:sub@fig_fixed_transmission_dimension_LQR, when sensor 1’s compression dimension increases, the LQR cost and state estimation error first decrease and then increase. The best result occurs at the allocation (4,2). As sensor 1 (the low-noise sensor) provides more reliable observations, increasing its compression dimension improves performance until lossless compression is effectively achieved at dimension 4. Beyond that, allocating additional resources to sensor 1 (e.g., (5,1)) reduces the transmission dimension available to sensor 2, discarding potentially valuable complementary information. Meanwhile, Fig. 6LABEL:sub@fig_fixed_transmission_dimension_MSE shows the reconstruction error of sensors 1 and 2 across allocation strategies. As sensor 1’s compression dimension increases, its reconstruction MSE decreases consistently. For sensor 2, as the compression dimension of sensor 1 increases, the MSE first decreases and then increases. This rise is due to two factors: initially, when the compression dimension of sensor 1 is low but that of sensor 2 is high, the LQR system deviates from its stable state, causing the sensor observations to fall outside the training distribution of AE and degrading the reconstruction quality. Later, as the compression dimension of sensor 1 increases, sensor 2 suffers from lossy compression at a lower compression dimension, and the reconstruction error increases again.
@@
Figure 7: Performance of the multi-sensor system with varying resource allocation and transmission dimensions.